# On the Fujita Phenomenon for a Forced Spatio-Temporal Fractional Diffusion Equation

### v1

@@ Abstract.
We study the Cauchy problem for a semilinear fractional diffusion equation with a time-dependent forcing term:
@@
with parameters $\alpha,\mathsf{s}\in(0,1)$, $\sigma>-\alpha$, and a continuous function $\mathbf{w}$. The operator $\partial_{t}^{\alpha}$ denotes the Caputo fractional derivative. Our main contributions are threefold. First, we prove the local existence of mild solutions and demonstrate a finite-time blow-up in the subcritical regime, provided $\int_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx>0$. Second, for the supercritical case with $\alpha<\sigma<0$, we establish the global existence for sufficiently small initial data and a forcing term, and we determine the critical exponent to be
@@
Finally, for this supercritical range, we prove a more robust global existence result under assumptions requiring only local smallness and controlled growth of the data.
@@
Fractional diffusion equation, local/global existence, blow-up, Fujita exponent.
@@ 1. Introduction and main results
In this paper, we study the Cauchy problem for a spatio-temporal fractional diffusion equation with a forcing term:
$\begin{cases}\partial_{t}^{\alpha}u+(-\Delta)^{\mathsf{s}}u=|u|^{p}+t^{\sigma}\mathbf{w}(x)&\text{in }\mathbb{R}^{N}\times(0,\infty),\\ u(x,0)=u_{0}(x)&\text{in }\mathbb{R}^{N},\end{cases}$ (1.1)
@@
Equations that are fully nonlocal, featuring both fractional time derivatives and fractional spatial operators, provide an effective framework for modeling phenomena with long-range interactions and memory effects. Such models naturally occur in a variety of physical settings, including plasma physics and turbulent transport—where anomalous diffusion arises from long-range correlations in space and time [9, 10]—geophysical flows and atmospheric dynamics, where memory and large-scale interactions play a fundamental role [34], as well as in financial mathematics for asset price dynamics incorporating long-term memory and jump processes [8].
From the analytical perspective, a central theme is the study of the critical exponent that delineates global existence from finite-time blow-up. The classical starting point is the seminal work of Fujita [17, 18] on the semilinear heat equation (corresponding to $\alpha=1$, $\mathsf{s}=1$ and $\mathbf{w}\equiv 0$ in our setting). Fujita proved that for $1<p<1+\frac{2}{N}$ and nontrivial initial data $u_{0}\not\equiv 0$, every solution blows up in finite time, while for $p>1+\frac{2}{N}$ global solutions exist for sufficiently small initial data. The critical value $p_{F}=1+\frac{2}{N}$ was later also shown to lie in the blow-up regime [22, 27]. These foundational results are thoroughly reviewed in Levine’s survey [13].
The Fujita phenomenon has been widely extended beyond the classical setting. Zhang and Sun [39] studied Problem (1.1) without a forcing term ($\mathbf{w}=0$) in the case $\mathsf{s}=1$ and $0<\alpha<1$, and established an analogous critical exponent $p_{F}=1+\frac{2}{N}$. For the fully fractional situation, Kirane et al. [21] analyzed the more general equation
@@
For the unforced version of our main problem (1.1) (that is, $\mathbf{w}=0$), the critical exponent can be inferred from the scaling properties of the equation. If $u(t,x)$ is a solution, then for any $\lambda>0$ the rescaled function
@@
which reduces to the classical value when $\alpha=\mathsf{s}=1$; see [16].
It is, however, well understood that critical exponents are highly sensitive to the structure of the problem. The addition of a forcing term may substantially alter the threshold between global existence and blow-up. For the classical heat equation, for example, forcing terms can shift the Fujita exponent [5]. Likewise, geometric effects may change the critical value, as demonstrated by Bandle and Levine [4] for conical domains.
Motivated by these considerations, and drawing on the general criticality framework developed in [19], we examine how the forcing term $t^{\sigma}\mathbf{w}(x)$ affects the Fujita exponent in the fully fractional setting of equation (1.1). Our objective is to determine whether, and in what manner, the critical exponent
$p_{F}=1+\frac{2\alpha}{N\alpha+2(1-\alpha)}$
obtained for the homogeneous case is modified in the presence of an external source. Our main result establishes a sharp critical exponent for the forced equation, precisely characterizing the transition between global existence and finite-time blow-up.
We begin by establishing local well-posedness in the Lebesgue space $L^{q}$ under appropriate assumptions.
@@
    The special case $\alpha=1$ with $\sigma=0$ was recently analyzed in [29]. A broader framework for the case $\alpha=1$ is developed in [7].
@@
  4. (iv)
    For related contributions in the setting $\alpha=\mathsf{s}=1$, we refer the reader to [2, 5, 23, 33]. In particular, when $\sigma\in(-1,0)$, one has
    $p^{*}=\frac{N-2\sigma}{\,N-2\sigma-2\,}$,
    which is consistent with [23, Theorem 1.1, (1.8)].
  5. (v)
    The special case $\alpha=1$ with integer $\mathsf{s}\geq 1$ was previously analyzed in [32].
  6. (vi)
    Our results show that the presence of the external forcing term $t^{\sigma}\mathbf{w}(x)$ accelerates finite-time blow-up compared to the dynamics observed in the unforced case.
  7. (vii)
    More specifically, the analysis indicates that the space-time dependent forcing term counteracts the regularizing effects of the spatio-temporal diffusion governed by the fractional operators $\partial_{t}^{\alpha}$ and $(-\Delta)^{\mathsf{s}}$.
  8. (viii)
  9. (ix)
@@
When a forcing term depends on both time and space, a weaker criterion for the global existence of $L^{q}$-solutions can be formulated by imposing smallness conditions on the local averages of the initial data $u_{0}$ and the forcing term $\mathbf{w}$ over balls of suitable radius.
@@
Outline of the paper. The paper is organized as follows. Section 2 collects preliminary material. We begin by reviewing essential concepts related to fractional derivatives and then introduce the fundamental kernels used to construct mild solutions via the Duhamel formula. The key properties of these kernels—self-similarity, integrability, and regularity—form the basis of our analysis of the existence and qualitative behavior of solutions. This section also establishes several technical lemmas that will serve as crucial tools for the subsequent estimates. In Section 3, we provide the proofs of the main results. We first establish a local existence result as stated in Theorem 1.1. Next, we address the nonexistence of global solutions by proving Theorem 1.2. We then prove the global existence result in Theorem 1.3, which relies on the smallness of the initial data $u_{0}$ in the critical norm and the smallness of the forcing term $\mathbf{w}$ in a suitable $L^{r_{c}}$ norm. Finally, we present Theorem 1.4, establishing global existence under significantly weaker local smallness conditions. Unlike Theorem 1.3, this result allows $u_{0}$ and $\mathbf{w}$ to be large in a global norm, provided they satisfy appropriate local smallness and decay assumptions.
@@ 2. Useful tools & Auxiliary results
We now state the following integration-by-parts formula holds for the Caputo derivative:
Lemma 2.1 (Integration by parts).
Let $\partial_{t}^{\alpha}u\in L^{1}(0,T)$, $v\in C([0,T])$, and $v(T)=0$. Then,
$\int\limits_{0}^{T}v(t)\partial_{t}^{\alpha}u(t),dt=\int\limits_{0}^{T}\left(u(t)-u(0)\right)\mathscr{D}_{t|T}^{\alpha}v(t)dt,$ (2.5)
where $\mathscr{D}_{t|T}^{\alpha}$ is the right-sided Riemann-Liouville fractional derivative of order $\alpha$, defined by
$\mathscr{D}_{t|T}^{\alpha}v(t)=\frac{1}{\Gamma(1-\alpha)}\frac{d}{dt}\int\limits_{t}^{T}(t-s)^{-\alpha}v(s)ds.$ (2.6)
@@
Using the definition of the Caputo derivative, we write
$\int\limits_{0}^{T}v(t)\,\partial_{t}^{\alpha}u(t)\,dt=\frac{1}{\Gamma(1-\alpha)}\int\limits_{0}^{T}v(t)\left(\int\limits_{0}^{t}(t-s)^{-\alpha}u^{\prime}(s)\,ds\right)dt.$
@@
We now integrate by parts, noting that $\frac{d}{ds}I^{1-\alpha}_{s|T}v(s)=-\mathscr{D}_{s|T}^{\alpha}v(s)$, and use the fact that $v(T)=0$:
$\int\limits_{0}^{T}u^{\prime}(s)\,I^{1-\alpha}_{s|T}v(s)\,ds=-\int\limits_{0}^{T}u(s)\,\mathscr{D}_{s|T}^{\alpha}v(s)\,ds+\left[u(s)I^{1-\alpha}_{s|T}v(s)\right]_{0}^{T}.$
Since $I^{1-\alpha}_{T|T}v(T)=0$, we obtain
$\int\limits_{0}^{T}v(t)\,\partial_{t}^{\alpha}u(t)\,dt=-\int\limits_{0}^{T}u(t)\,\mathscr{D}_{t|T}^{\alpha}v(t)\,dt+u(0)I^{1-\alpha}_{0|T}v(0).$
It remains to compute the integral of the right-sided Riemann–Liouville derivative:
$\int\limits_{0}^{T}\mathscr{D}_{t|T}^{\alpha}v(t)\,dt=\frac{1}{\Gamma(1-\alpha)}\int\limits_{0}^{T}\frac{d}{dt}\left(\int\limits_{t}^{T}(s-t)^{-\alpha}v(s)\,ds\right)dt.$
Using Fubini again, we obtain
$\frac{1}{\Gamma(1-\alpha)}\int\limits_{0}^{T}s^{-\alpha}v(s)\,ds=-I^{1-\alpha}_{0|T}v(0),$
and thus
$u(0)\int\limits_{0}^{T}\mathscr{D}_{t|T}^{\alpha}v(t)\,dt=-u(0)I^{1-\alpha}_{0|T}v(0).$
Substituting back, we find
$\int\limits_{0}^{T}v(t)\,\partial_{t}^{\alpha}u(t)\,dt=\int\limits_{0}^{T}(u(t)-u(0))\mathscr{D}_{t|T}^{\alpha}v(t)\,dt,$
which proves the result. ∎
@@ 2.1. Fundamental solution and Duhamel’s formula
$\partial_{t}^{\alpha}u+(-\Delta)^{\mathsf{s}}u=0,$
while $Y=Z$ when $\alpha=1$, and $Y=\partial_{t}^{\alpha}Z$ for $\alpha\in(0,1)$; see [11, 24].
@@
    For $t\geq t_{0}>0$,
@@
    where the constant depending only on $\alpha,\mathsf{s},N$, and q.
@@
It is straightforward to verify that
@@
where $p_{c}$ is given in (1.3). Let $u$ be a maximal $L^{q}$-solution of (1.1) corresponding to initial data $u_{0}\in L^{q}(\mathbb{R}^{N})$, and assume $\mathbf{w}\in L^{q}(\mathbb{R}^{N})$.
Then there exists a constant $\mathscr{M}>0$, depending only on $p,q,\alpha,\mathsf{s},N$, and $\sigma$, such that if
@@
then the following a apriori bound holds:

### v2

@@ Abstract.
We investigate the Cauchy problem for a semilinear spatio–temporal fractional diffusion equation with a time-dependent forcing term:
@@
where $\alpha,\mathsf{s}\in(0,1)$, $\sigma>-\alpha$, and $\mathbf{w}$ is a given continuous function. Here $\partial_{t}^{\alpha}$ denotes the Caputo fractional derivative. Our main results are threefold. First, we establish local-in-time existence of mild solutions and prove finite-time blow-up in the subcritical regime, under the positivity condition
$\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx>0.$
Second, in the supercritical case $\alpha<\sigma<0$, we prove the global existence of solutions for sufficiently small initial data and forcing term, and we identify the corresponding critical exponent as
@@
Finally, within this supercritical range, we obtain a more robust global existence result under weaker assumptions that require only local smallness and controlled growth of the data. To the best of our knowledge, a sharp Fujita-type threshold for fully spatio-temporal fractional diffusion equations with time-growing external forcing has not been previously established.
@@
Time–fractional diffusion equation; fractional Laplacian; Fujita phenomenon; critical exponent; finite-time blow-up; global existence; external forcing.
@@ 1. Introduction and main results
In this paper, we investigate the Cauchy problem for a spatio-temporal fractional diffusion equation subject to an external forcing term:
$\begin{cases}\partial_{t}^{\alpha}u+(-\Delta)^{\mathsf{s}}u=|u|^{p}+t^{\sigma}\mathbf{w}(x)&\text{in }(0,\infty)\times\mathbb{R}^{N},\\ u(0,x)=u_{0}(x)\geq 0&\text{in }\mathbb{R}^{N},\end{cases}$ (1.1)
@@
The framework of fully nonlocal equations, which combines fractional time derivatives with fractional spatial operators, offers a powerful approach for capturing phenomena characterized by long-range interactions and memory effects. Such models naturally occur in a variety of physical settings, including plasma physics and turbulent transport where anomalous diffusion arises from long-range correlations in space and time [9, 10] geophysical flows and atmospheric dynamics, where memory and large scale interactions play a fundamental role [34], as well as in financial mathematics for asset price dynamics incorporating long-term memory and jump processes [8].
From an analytical point of view, identifying the critical exponent that separates global-in-time existence from finite-time blow-up is a central theme. The classic starting point is Fujita’s seminal work [17, 18] on the semi-linear heat equation, which corresponds in our framework to the case $\alpha=1$, $\mathsf{s}=1$ and $\mathbf{w}\equiv 0$. He established that for $1<p<1+\frac{2}{N}$, any solution resulting from a non-trivial initial condition $u_{0}\not\equiv 0$ blows up in finite time, whereas for $p>1+\frac{2}{N}$, global solutions exist if the initial condition is sufficiently small. The critical value $p_{F}=1+\frac{2}{N}$ was subsequently incorporated into the blow-up regime by Hayakawa [22, 27]. These fundamental results are detailed in Levine’s summary [13].
Subsequently, the research then focused on specific settings. For $\alpha=1$ and $\sigma=0$, the problem was analyzed in [29], while a more general framework with $\alpha=1$ was later developed in [7]. Related nonlocal configurations involving different fractional diffusion operators were also studied in [6].
In the classical setting corresponding to $\alpha=\mathsf{s}=1$, there are many related works in the literature, notably [2, 5, 23, 33]. In particular, for $\sigma\in(-1,0)$, the critical exponent admits the explicit expression
$p_{F}=\frac{N-2\sigma}{\,N-2\sigma-2\,}$,
which coincides with the result [23, Theorem 1.1, (1.8)]. The case $\alpha=1$ with an integer spatial order $\mathsf{s}\geq 1$ has also been studied in detail in [32].
The Fujita phenomenon was later generalized beyond the classical framework. For Problem (1.1) without a forcing term, with $\mathsf{s}=1$ and $0<\alpha<1$, Zhang and Sun [39] determined the analogous critical exponent $p_{F}=1+\frac{2}{N}$. For the fully fractional situation, Kirane et al. [21] analyzed the more general equation
@@
For the unforced version of our main problem (1.1), the critical exponent can be inferred from the scaling properties of the equation. If $u(t,x)$ is a solution, then for any $\lambda>0$ the rescaled function
@@
which reduces to the classical value when $\alpha=\mathsf{s}=1$.
It is well known that critical exponents are highly sensitive to the underlying structure of the equation. In particular, the presence of additional mechanisms may significantly modify the threshold separating global-in-time existence from finite-time blow-up. For the classical heat equation, for instance, it has been shown that forcing terms can shift the Fujita exponent [5]. Likewise, geometric effects may alter the critical value, as demonstrated by Bandle and Levine [4] in the case of conical domains.
A central contribution of the present work is the precise characterization of the impact of an external forcing term on the qualitative behavior of solutions. Our analysis relies on the general criticality framework developed in [21]. In this setting, we investigate how the critical blow-up exponent associated with the unforced problem,
$p_{F}=1+\frac{2\mathsf{s}\alpha}{N\alpha+2\mathsf{s}(1-\alpha)},$
is modified by the inclusion of an external source.
Our main results identify a sharp critical exponent for the forced equation, which accurately delineates the transition between global-in-time existence and finite-time blow-up. In particular, they show that the presence of a forcing term accelerates the blow-up mechanism when compared to the unforced case, in the sense that the range of exponents $p$ for which solutions necessarily blow up in finite time becomes strictly larger. More precisely, space- and time-dependent forcing effects tend to counterbalance the regularizing influence of fractional diffusion, as governed by the operators $\partial_{t}^{\alpha}$ and $(-\Delta)^{\mathsf{s}}$.
From a modeling perspective, the incorporation of a time-dependent forcing term of the form $t^{\sigma}\mathbf{w}(x)$ is well motivated in physical systems exhibiting anomalous diffusion and long-memory effects. Time-fractional diffusion equations involving the Caputo derivative are widely employed to describe subdiffusive transport in disordered media, plasma physics, and turbulent transport, where nonlocal temporal effects capture the persistent influence of past states [9, 10, 14, 25]. In such contexts, external sources typically act continuously in time rather than instantaneously, giving rise to forcing mechanisms whose intensity evolves according to a power law. Closely related phenomena also occur in geophysical flows and atmospheric dynamics, where sustained energy injection interacts with nonlocal diffusion processes [34]. The exponent $\sigma$ thus encodes the temporal balance between external forcing and fractional dissipation, and its interplay with the nonlinear reaction term can have a decisive impact on the qualitative behavior of solutions. To rigorously quantify the effect of external forcing on the classical Fujita threshold and on the global dynamics of the system, it is necessary to first establish a local well-posedness theory. To this end, we prove the existence and uniqueness of mild solutions in suitable Lebesgue spaces, which provides the analytical foundation for the subsequent study of finite-time blow-up and global-in-time existence. Accordingly, we begin our analysis by deriving a local well-posedness result in the Lebesgue space framework under appropriate assumptions on the data.
@@
When a forcing term depends on both time and space, a weaker criterion for the global existence of $L^{q}$-solutions can be formulated by imposing smallness conditions on the local averages of the initial data $u_{0}$ and the forcing term $\mathbf{w}$ over balls of suitable radius. The following result provides a global existence criterion based on local averages rather than norm smallness.
@@
Outline of the paper. The paper is organized as follows.
In Section 2, we collect the preliminary material needed throughout the paper. We begin by reviewing essential concepts related to fractional derivatives and then introduce the fundamental kernels used to construct mild solutions via the Duhamel formula. The key properties of these kernels—such as self-similarity, integrability, and regularity—form the basis of our analysis of the existence and qualitative behavior of solutions. This section also establishes several technical lemmas that serve as crucial tools for the estimates developed later.
Section 3 is devoted to the proofs of the main results. We first establish a local existence result, stated in Theorem 1.1. We then address the nonexistence of global solutions by proving Theorem 1.3. Next, we prove the global existence result in Theorem 1.5, which relies on the smallness of the initial data $u_{0}$ in the critical norm and the smallness of the forcing term $\mathbf{w}$ in a suitable $L^{r_{c}}$ norm. Finally, we present Theorem 1.7, establishing global existence under significantly weaker local smallness conditions. Unlike Theorem 1.5, this result allows $u_{0}$ and $\mathbf{w}$ to be large in a global norm, provided they satisfy appropriate local smallness and decay assumptions.
In Section 4, we complement the analytical results with numerical simulations of a one-dimensional fractional diffusion equation. The simulations are designed to illustrate the critical threshold predicted by Theorem 1.3 and to visualize the qualitative differences between subcritical and supercritical regimes. In particular, we compare the time evolution and spatial profiles of solutions for exponents below and above the critical value, highlighting the concentration mechanism leading to blow-up in the subcritical case and the bounded behavior observed in the supercritical regime. These computations do not constitute a proof, but they provide a concrete and intuitive illustration of the theoretical phenomena established in the paper.
@@ 2. Useful tools & Auxiliary results
We now present the following lemma concerning the fractional integration-by-parts formula for the Caputo derivative.
Lemma 2.1.
Let $\alpha\in(0,1)$, $u\in AC([0,T])$, and $v\in C^{1}([0,T])$ with $v(T)=0$. Then,
$\int\limits_{0}^{T}v(t)\,\partial_{t}^{\alpha}u(t)\,dt=\int\limits_{0}^{T}\big(u(t)-u(0)\big)\,\mathscr{D}_{t|T}^{\alpha}v(t)\,dt,$ (2.5)
where $\mathscr{D}_{t|T}^{\alpha}$ is the right-sided Riemann-Liouville fractional derivative defined by
$\mathscr{D}_{t|T}^{\alpha}v(t)=-\frac{1}{\Gamma(1-\alpha)}\frac{d}{dt}\int\limits_{t}^{T}(s-t)^{-\alpha}v(s)\,ds.$ (2.6)
@@
The result follows directly from the standard properties of the Caputo derivative and Riemann-Liouville fractional derivatives; see, for instance, [35, Section 2.4].
Recall that for $0<\alpha<1$, the Caputo derivative of order $\alpha$ is defined as Thus,
$\int_{0}^{T}v(t)\,\partial_{t}^{\alpha}u(t)\,dt=\frac{1}{\Gamma(1-\alpha)}\int_{0}^{T}v(t)\left(\int_{0}^{t}(t-s)^{-\alpha}u^{\prime}(s)\,ds\right)dt.$
@@
We integrate by parts, we write
$\int_{0}^{T}u^{\prime}(s)\,I_{s|T}^{1-\alpha}v(s)\,ds=\Big[u(s)\,I_{s|T}^{1-\alpha}v(s)\Big]_{s=0}^{s=T}-\int_{0}^{T}u(s)\,\frac{d}{ds}I_{s|T}^{1-\alpha}v(s)\,ds.$
Now compute the derivative of the fractional integral. By the definition of the right-sided Riemann-Liouville derivative (2.6), we have
$\frac{d}{ds}I_{s|T}^{1-\alpha}v(s)=\frac{1}{\Gamma(1-\alpha)}\frac{d}{ds}\int_{s}^{T}(t-s)^{-\alpha}v(t)\,dt=-\mathscr{D}_{s|T}^{\alpha}v(s).$
Evaluate the boundary term. Since $I_{T|T}^{1-\alpha}v(T)=0$ (the integral is over an interval of zero length), we obtain
$\Big[u(s)\,I_{s|T}^{1-\alpha}v(s)\Big]_{s=0}^{s=T}=u(T)\cdot 0-u(0)\,I_{0|T}^{1-\alpha}v(0)=-u(0)\,I_{0|T}^{1-\alpha}v(0).$
Therefore,
$\int_{0}^{T}v(t)\,\partial_{t}^{\alpha}u(t)\,dt=-\int_{0}^{T}u(s)\,\frac{d}{ds}I_{s|T}^{1-\alpha}v(s)\,ds-u(0)I_{0|T}^{1-\alpha}v(0)=\int_{0}^{T}u(s)\,\mathscr{D}_{s|T}^{\alpha}v(s)\,ds-u(0)I_{0|T}^{1-\alpha}v(0).$
It remains to express $I_{0|T}^{1-\alpha}v(0)$ in terms of $\mathscr{D}_{t|T}^{\alpha}v(t)$. Compute the integral of this derivative:
$\int_{0}^{T}\mathscr{D}_{t|T}^{\alpha}v(t)\,dt=-\frac{1}{\Gamma(1-\alpha)}\int_{0}^{T}\frac{d}{dt}\left(\int_{t}^{T}(s-t)^{-\alpha}v(s)\,ds\right)dt.$
By the Fundamental Theorem of Calculus,
$\frac{1}{\Gamma(1-\alpha)}\left[\int_{t}^{T}(s-t)^{-\alpha}v(s)\,ds\right]_{t=0}^{t=T}.$
For $t=T$, the integral $\int_{T}^{T}(s-T)^{-\alpha}v(s)\,ds=0$. For $t=0$, we obtain $\int_{0}^{T}s^{-\alpha}v(s)\,ds$. Thus,
$\int_{0}^{T}\mathscr{D}_{t|T}^{\alpha}v(t)\,dt=\frac{1}{\Gamma(1-\alpha)}\int_{0}^{T}s^{-\alpha}v(s)\,ds=I_{0|T}^{1-\alpha}v(0).$
We deduce that
$u(0)I_{0|T}^{1-\alpha}v(0)=-u(0)\int_{0}^{T}\mathscr{D}_{t|T}^{\alpha}v(t)\,dt.$
Substituting back into the previous expression yields
$\int_{0}^{T}v(t)\,\partial_{t}^{\alpha}u(t)\,dt=\int_{0}^{T}u(t)\,\mathscr{D}_{t|T}^{\alpha}v(t)\,dt-u(0)\int_{0}^{T}\mathscr{D}_{t|T}^{\alpha}v(t)\,dt=\int_{0}^{T}\big(u(t)-u(0)\big)\,\mathscr{D}_{t|T}^{\alpha}v(t)\,dt,$
which establishes the desired integration by parts formula. ∎
@@ 2.1. Fundamental solution and Duhamel’s formula
$\partial_{t}^{\alpha}u+(-\Delta)^{\mathsf{s}}u=0.$
For $\alpha\in(0,1)$, the associated kernel $Y$ which corresponds to the time fractional Duhamel term satisfies $Y(t,\cdot)\sim\partial_{t}^{\alpha}Z(t,\cdot)$ in the sense of distributions [11, 24].
@@
    For $t\geq t_{0}>0$, there exists $C=C(\alpha,\mathsf{s},N,q)>0$ such that
@@
A direct computation shows that
@@
where $p_{c}$ is given in (1.3). Let $u$ be a maximal $L^{q}$-solution of (1.1) corresponding to initial data $u_{0}\in L^{q}(\mathbb{R}^{N})$, and assume $\mathbf{w}\in L^{q}(\mathbb{R}^{N})$. Then there exists a constant $\mathscr{M}>0$, depending only on $p,q,\alpha,\mathsf{s},N$, and $\sigma$, such that if
@@
then the following an a priori bound: